OpenStudy (anonymous):

Need help please. with cherrys on top :D (x^3+x^2+2x-4÷(x-1) Anwsers: 1. x2 + 2x + 4 2. x2 + x - 4 3. x2 + 4 10x^3-8x^2-12x/2x Anwsers: 5x2 - 4x - 6 5x6 - 4x4 - 6x4 5x3 - 4x2 - 6x

OpenStudy (campbell_st):

(x -1) is a factor as f(1) = 0. Use some polynomial division to solve it. Alternatively take (x -1) and multiply by each answer and see which gives the original equation

OpenStudy (dumbcow):

For 1st one use synthetic division: the zero is 1 1| 1 1 2 -4 1 2 4 ------------ 1 2 4 0 The quotient is then : x^2 +2x +4 For the 2nd one just divide each term: 10x^3/2x = 5x^2 -8x^2/2x = -4x -12x/2x = -6 --> 5x^2 -4x -6